OpenStudy (fanduekisses):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (shamim):

u know n-th term of a geometri series is tn=ar^(n-1)

OpenStudy (shamim):

so second tetm t2=ar^(2-1)=-12 ar=-12

OpenStudy (shamim):

now 5 th term is t5=ar^(5-1)=ar^4=768

OpenStudy (shamim):

now devide those two equation

OpenStudy (fanduekisses):

-64

OpenStudy (shamim):

yes

OpenStudy (shamim):

r=?

OpenStudy (fanduekisses):

is -64?

OpenStudy (shamim):

r^3=-64 r=-4

OpenStudy (shamim):

again ar=-12 a(-4)=-12 a=?

OpenStudy (fanduekisses):

a=3

OpenStudy (shamim):

now nth term is tn=ar^(n-1)=?

OpenStudy (fanduekisses):

What I don't quite understand is why divide the two equations to get -64 and they how we get r=-4?